# Fortune: Formula-Driven Reinforcement Learning for Symbolic Table Reasoning in Language Models

### v2

@@ Appendix F Ablation Study and Upper-Bound Performance of Fortune++
TabAF [22] Upper Bound 80.13 94.02 - 82.07 - - -
@@
Fortune Upper Bound 77.35 96.49 72.01 79.91 54.82 88.93 43.26
@@
Fortune++ Upper Bound 93.62 99.51 91.02 95.01 71.29 98.06 61.16
@@
We conduct an ablation study of Fortune++ to investigate the complementary roles and effectiveness of textual and symbolic reasoning. Table 4 reveals several instructive patterns.
@@
Combining textual and symbolic reasoning yields the most robust performance. The balanced sampling strategy used by Fortune++ (five textual and five formula candidates) consistently outperforms both the pure-text and pure-formula variants across all benchmarks. This confirms that textual and symbolic reasoning address complementary error modes. Disabling either modality leads to substantial accuracy degradation, with drops of up to 18 percentage points (e.g., –18 pp on FinQA for text-only, –13 pp on AIT-QA for formula-only).
@@
Text and formula reasoning each excel in different scenarios. Textual reasoning performs better on simpler table QA tasks, such as TabFact and AIT-QA, where natural language understanding and logical inference dominate. In contrast, formula-based reasoning excels on arithmetic-heavy or structured computation tasks like FinQA and MultiHiertt, where symbolic execution is crucial for deriving the correct answer. This division reinforces the importance of integrating both modalities for general-purpose table understanding.
@@
Reinforcement learning enhances symbolic reasoning beyond supervised fine-tuning. Compared with TabAF—which uses the same backbone but is trained only with SFT—Fortune’s RL variant achieves substantially stronger formula-only performance (e.g., 82.41% vs. 67.54% on TabFact with 5 Formula). This suggests that reinforcement learning encourages the model to explore more reliable and executable reasoning paths, ultimately improving symbolic program quality.
@@
Many correct answers are lost due to naive majority voting. The upper-bound rows show that Fortune++ frequently generates correct answers that are not selected by simple majority vote. The discrepancy between upper-bound and actual performance reaches 19 pp on MultiHiertt and 17 pp on TableBench, indicating considerable headroom for improved candidate selection through confidence-based aggregation or smarter reranking mechanisms.
@@
Complex, low-resource benchmarks remain the greatest challenge. The largest performance gaps appear on structurally complex and low-resource datasets such as MultiHiertt and TableBench. These results highlight the limitations of current voting and reasoning mechanisms and point to future directions including symbolic planner integration, adaptive sampling, and confidence-calibrated answer selection.
Appendix G Comparative Analysis of Formula Tuning and Other Symbolic Table Reasoning Methods
(a) Zero-Shot setting
(b) Reinforcement Learning setting
Figure 3: Comparison of performance across different symbolic reasoning methods under Zero-Shot and Reinforcement Learning (RL) settings. Values in the figure indicate accuracy (%).
Figure 3 and Table 5 compare three symbolic reasoning paradigms—SQL queries, Python snippets, and spreadsheet formulas—under both zero-shot and reinforcement learning (RL) settings. To ensure a fair comparison, we train and evaluate these symbolic tools only on datasets with flat, relational tables: WikiTQ, TabFact, and FinQA for training, and all three plus TableBench for evaluation. This restriction is necessary because SQL operates solely on flat tables, while Python/Pandas are also usually worked for flat tabular structures. Several patterns emerge from the comparison.
Method WikiTQ TabFact FinQA TableBench Overall
Zero-Shot
   SQL 17.07 21.15 1.51 7.64 11.84
   Python 30.96 60.39 34.87 16.33 35.64
   Formula 38.96 52.52 34.44 26.05 37.99
Reinforcement Learning (RL)
   SQL 67.58 83.94 38.79 32.09 55.60
   Python 70.46 84.42 65.85 35.26 64.00
   Formula 70.67 84.50 65.89 35.84 64.23
@@ Appendix H Impact Analysis of the Thinking Process in Formula Tuning
In the zero-shot setting, reasoning may hurts performance. We observe that adding a reasoning trace before the formula generally leads to lower accuracy (e.g., 39.24% → 36.13% overall). This is likely because such thought-first-then-formula generation patterns are underrepresented in pretraining corpora. As a result, models tend to produce unnatural or error-prone reasoning steps, which negatively affect the final output. The limitations of chain-of-thought reasoning in certain scenarios have also been discussed in recent work [82, 83].
@@ Appendix I Statistical Analysis of Generated Formulas in Symbolic Table Reasoning
Spreadsheet formulas are Turing complete [23], but in practice, only a core subset of functions is commonly used. To guide the language model toward effective formula generation and support future research, we identify a set of representative operators in Table 11. This table lists each operator’s name, symbol, definition, and a representative example—primarily covering basic arithmetic and aggregation operations. Acknowledging the model’s limited formula knowledge at the beginning of training, we explicitly introduce these operators during prompting, while still allowing the model to use any formula supported by our execution engine.
@@
Figures 6, 7, 8, 9, and 10 illustrate the prompts used in our experiments across zero-shot inference, supervised fine-tuning (SFT), reinforcement learning (RL), and evaluation.
Pre-text and Post-text refer to optional unstructured context surrounding the table, such as the data description format used in FinQA [63]. formula operator instruction represents the textual representations and usage guidelines of the representative spreadsheet formula operators in symbolic table reasoning, as detailed in Appendix K.
@@
Here is the LAST PAGE.

### v3

@@ Appendix G Ablation Study and Upper-Bound Performance of Formula-R1
TabAF (Wang et al., 2025) Upper Bound 80.13 94.02 - 82.07 - - -
@@
Formula-R1 Upper Bound 77.35 96.49 72.01 79.91 54.82 88.93 43.26
@@
Formula-R1-Plus Upper Bound 93.62 99.51 91.02 95.01 71.29 98.06 61.16
@@
We conduct an ablation study of Formula-R1-Plus to investigate the complementary roles and effectiveness of textual and symbolic reasoning. Table 6 reveals several notable patterns.
@@
Combining textual and symbolic reasoning yields the most robust performance. The balanced sampling strategy adopted by Formula-R1-Plus (five textual and five formula-based candidates) consistently outperforms both the pure-text and pure-formula variants across all benchmarks. This result suggests that textual and symbolic reasoning address complementary error modes. Disabling either modality leads to substantial accuracy drops, with degradations of up to 18 percentage points (e.g., 18 points on FinQA for text-only and 13 points on AIT-QA for formula-only).
@@
Textual and formula-based reasoning excel in different scenarios. Textual reasoning performs better on relatively simpler table QA tasks, such as TabFact and AIT-QA, where natural language understanding and logical inference play a dominant role. In contrast, formula-based reasoning is more effective on arithmetic-intensive or structurally complex tasks such as FinQA and MultiHiertt, where symbolic execution is essential for deriving the correct answer. This contrast further highlights the importance of integrating both modalities for general-purpose table understanding.
@@
Reinforcement learning enhances symbolic reasoning beyond supervised fine-tuning. Compared with TabAF, which uses the same backbone but is trained solely with SFT, the RL-based variant of Formula-R1 achieves substantially stronger formula-only performance (e.g., 82.41% vs. 67.54% on TabFact with 5 Formula). This result suggests that reinforcement learning encourages the model to explore more reliable and executable reasoning paths, ultimately improving the quality of symbolic programs.
@@
A substantial number of correct answers are lost under naive majority voting. The upper-bound rows show that Formula-R1-Plus frequently generates correct answers that are not selected by simple majority voting. The gap between the upper bound and actual performance reaches 19 points on MultiHiertt and 17 points on TableBench, indicating considerable headroom for improved candidate selection through confidence-based aggregation or more effective reranking strategies.
@@
Structurally complex and low-resource benchmarks remain the most challenging. The largest performance gaps appear on datasets such as MultiHiertt and TableBench, which are both structurally complex and relatively low-resource. These results highlight the limitations of current voting and reasoning mechanisms, and point to promising future directions such as symbolic planner integration, adaptive sampling, and confidence-calibrated answer selection.
@@ Appendix H Impact Analysis of the Thinking Process in Formula Tuning
In the zero-shot setting, reasoning may hurts performance. We observe that adding a reasoning trace before the formula generally leads to lower accuracy (e.g., 39.24% → 36.13% overall). This is likely because such thought-first-then-formula generation patterns are underrepresented in pretraining corpora. As a result, models tend to produce unnatural or error-prone reasoning steps, which negatively affect the final output. The limitations of chain-of-thought reasoning in certain scenarios have also been discussed in recent work (Sprague et al., 2025; Liu et al., 2024).
@@ Appendix I Statistical Analysis of Generated Formulas in Symbolic Table Reasoning
Spreadsheet formulas are Turing complete (Smalley, 2023), but in practice, only a core subset of functions is commonly used. To guide the language model toward effective formula generation and support future research, we identify a set of representative operators in Table 12. This table lists each operator’s name, symbol, definition, and a representative example—primarily covering basic arithmetic and aggregation operations. Acknowledging the model’s limited formula knowledge at the beginning of training, we explicitly introduce these operators during prompting, while still allowing the model to use any formula supported by our execution engine.
@@
Figures 5, 6, 7, 8, and 9 illustrate the prompts used in our experiments across zero-shot inference, supervised fine-tuning (SFT), reinforcement learning (RL), and evaluation.
Pre-text and Post-text refer to optional unstructured context surrounding the table, such as the data description format used in FinQA (Chen et al., 2021b). formula operator instruction represents the textual representations and usage guidelines of the representative spreadsheet formula operators in symbolic table reasoning, as detailed in Appendix K.
@@
Appendix N The Use of Large Language Models (LLMs)
In this paper, we use an LLM (GPT-4o) as the source for distilling the supervision data used in the cold-start SFT stage. In addition, LLMs were used to assist with manuscript writing and polishing.